OpenStudy (anonymous):

Evaluate the following expression. (xy) ^0

OpenStudy (anonymous):

\[(xy)^0\]

OpenStudy (khaakenajaf):

x^0 y^0=1 *1=1

OpenStudy (anonymous):

anything to the power of 0 is always 1